# Reinforcement Unlearning

### v2

@@ 6. Conclusion
This paper presents a pioneering research area, termed reinforcement unlearning, which addresses the crucial need to protect the privacy of environment owners by enabling an agent to unlearn entire environments. We propose two distinct reinforcement unlearning methods: decremental RL-based and environment poisoning-based approaches. These methods are designed to be adaptable to different situations and provide effective mechanisms for unlearning. Also, we introduce a novel concept termed “environment inference” to evaluate the outcomes of the unlearning process. This evaluation framework allows us to assess the efficacy of our unlearning methods and gauge the level of privacy protection achieved in reinforcement learning-driven critical applications.
@@
Additional experimental results, along with details regarding the model architecture, are provided in this appendix.
1. Model Architecture
In the grid world and aircraft landing settings, we employ a fully-connected neural network as the model for our reinforcement learning agent. The neural network architecture consists of an input layer, two hidden layers, and an output layer. The input layer takes a 10-dimensional vector as input, representing the relevant features of the environment. The output layer generates a 4-dimensional vector, representing the probability distribution over the four possible actions: up, down, left, and right. The network architecture incorporates two hidden layers to facilitate learning and representation of complex patterns. The first hidden layer comprises 64 neurons, while the second one consists of 32 neurons.
@@ 1. Model Architecture
2. Overall Performance
Refer to caption (a) The average number of steps before and after unlearning, where Environment 1 is the unlearning environment Refer to caption (b) The average rewards received by the agent before and after unlearning, where Environment 1 is the unlearning environment
Figure 17. The decremental RL-based method in Grid World
Refer to caption (a) The average number of steps before and after unlearning, where Environment 1 is the unlearning environment Refer to caption (b) The average rewards received by the agent before and after unlearning, where Environment 1 is the unlearning environment
Figure 18. The poisoning-based method in Grid World
However, upon closer comparison between Figures 17 and 18, we can observe slight differences in the performance of the two methods in some remaining environments, such as Environments 18 and 19. In these environments, the poisoning-based method maintains almost unchanged steps and rewards between the pre-unlearning and post-unlearning stages, while the decremental reinforcement learning-based method does not achieve this. This result suggests that the decremental reinforcement learning-based method can potentially suffer from the over-unlearning issue to some extent, while the poisoning-based method demonstrates its ability to overcome this issue and retain better performance in the remaining environments after unlearning. These findings highlight the different characteristics and strengths of the two unlearning methods.
Figures 19, 20 and 21 demonstrate the overall performance of the decremental RL-based method in the context of aircraft landing, virtual home and maze explorer, respectively. In all the three scenarios, the agent’s behavior exhibits a notable increase in steps taken and a significant decrease in rewards achieved after the unlearning process. The reason for these trends in the three scenarios is rooted in the fundamental nature of reinforcement unlearning. The unlearning process seeks to selectively modify the agent’s behavior to forget specific environments or aspects of its learning history. Thus, the agent must re-explore and adapt to new circumstances, leading to fluctuations in its performance.
Refer to caption (a) The average number of steps before and after unlearning, where Environment 1 is the unlearning environment Refer to caption (b) The average rewards received by the agent before and after unlearning, where Environment 1 is the unlearning environment
Figure 19. The decremental RL-based method in Aircraft Landing
Refer to caption (a) The average number of steps before and after unlearning, where Environment 1 is the unlearning environment Refer to caption (b) The average rewards received by the agent before and after unlearning, where Environment 1 is the unlearning environment
Figure 20. The decremental RL-based method in Virtual Home
Refer to caption (a) The average number of steps before and after unlearning, where Environment 1 is the unlearning environment Refer to caption (b) The average rewards received by the agent before and after unlearning, where Environment 1 is the unlearning environment
Figure 21. The decremental RL-based method in Maze Explorer
Figures 22, 23 and 24 provide a comprehensive view of the poisoning-based method’s performance in the aircraft landing, virtual home and maze explorer settings, respectively. Remarkably, the performance trend in all the three scenarios is similar to that of the decremental RL-based method. This observation reinforces the effectiveness of the poisoning-based approach in reinforcement unlearning, as it consistently achieves the objective of degrading the agent’s performance in the targeted unlearning environment while preserving its capabilities in other environments. The consistent performance trend across different settings shows the method’s versatility and potential applicability in various RL scenarios.
Refer to caption (a) The average number of steps before and after unlearning, where Environment 1 is the unlearning environment Refer to caption (b) The average rewards received by the agent before and after unlearning, where Environment 1 is the unlearning environment
Figure 22. The poisoning-based method in Aircraft Landing
Refer to caption (a) The average number of steps before and after unlearning, where Environment 1 is the unlearning environment Refer to caption (b) The average rewards received by the agent before and after unlearning, where Environment 1 is the unlearning environment
Figure 23. The poisoning-based method in Virtual Home
Refer to caption (a) The average number of steps before and after unlearning, where Environment 1 is the unlearning environment Refer to caption (b) The average rewards received by the agent before and after unlearning, where Environment 1 is the unlearning environment
Figure 24. The poisoning-based method in Maze Explorer
3. Hyperparameter Study
Impact of Environment Size. In the aircraft landing setting (Figure 25), the discrepancy in rewards between the pre-unlearning and post-unlearning stages becomes more pronounced for both the decremental RL-based and poisoning-based methods. Interestingly, this discrepancy in the aircraft landing setting is even larger compared to that observed in the grid world setting. The reason behind this distinction can be attributed to the inherent differences in the features and complexities of the two settings. In the grid world setting, the environment primarily consists of discrete and structured grids, with agent navigating through straightforward paths. The relatively limited environment complexity in the grid world setting allows the unlearning methods to efficiently adapt the agent’s behavior and modify its policy, resulting in significant yet manageable changes in rewards. On the other hand, the aircraft landing setting is considerably more intricate and continuous, involving multiple variables and parameters governing the aircraft’s landing procedures. As the size of the environment increases, the number of possible landing trajectories and configurations expands exponentially. This complexity poses a greater challenge to the unlearning methods, requiring them to navigate a more vast policy space.
Refer to caption (a) Average steps of the decremental RL-based method Refer to caption (b) Rewards of the decremental RL-based method Refer to caption (c) Average steps of the poisoning-based method Refer to caption (d) Rewards of the poisoning-based method
Figure 25. The drecremental RL-based and poisoning-based methods in Aircraft Landing with different size
Impact of Environment Complexity. The complexity of an environment can be characterized by the presence and arrangement of obstacles within it. By modifying the complexity of the environment, we can assess the adaptability of the proposed methods. In the grid world setting, we examine the impact of increasing the environment complexity by introducing more obstacles, with the environment size maintained at $10\times 10$. Figures 26(a) and 26(b) show the outcomes of the decremental RL-based method.
Refer to caption (a) Steps of decremental RL-based Refer to caption (b) Rewards of decremental RL-based Refer to caption (c) Steps of poisoning-based Refer to caption (d) Rewards of poisoning-based
Figure 26. The decremental RL-based and poisoning-based methods in Grid World with different complexity
However, when employing the poisoning-based method (depicted in Figures 26(c) and 26(d)), an interesting observation emerges. Unlike the decremental RL-based method, the difference in both steps and rewards between the pre-unlearning and post-unlearning stages remains relatively stable even as the number of obstacles increases. The reason behind this intriguing behavior lies in the nature of the poisoning-based approach. When we introduce additional obstacles to the environment, the poisoning-based method operates differently compared to the decremental RL-based approach. Instead of modifying the agent’s learned policy gradually, the poisoning-based method incorporates an element of targeted perturbation. As the number of obstacles increases, the poisoning-based method strategically poisons the agent’s policy by introducing deceptive information during the unlearning process. This targeted perturbation causes the agent’s behavior to deviate from the optimal path more significantly, leading to relatively constant differences in both step count and reward between the pre-unlearning and post-unlearning phases.
Secondly, the poisoning-based method’s targeted perturbations are designed to cause deliberate deviations from the optimal path. As a result, the agent’s policy becomes consistently misled in the presence of additional obstacles, leading to a stable performance difference between the pre-unlearning and post-unlearning states.
In Figure 27, the similar results observed in the context of aircraft landing corroborate the findings in the grid world setting. This outcome can be attributed to the analogous reasons previously identified in the grid world scenario.
Refer to caption (a) Average steps of the decremental RL-based method Refer to caption (b) Rewards of the decremental RL-based method Refer to caption (c) Average steps of the poisoning-based method Refer to caption (d) Rewards of the poisoning-based method
Figure 27. The drecremental RL-based and poisoning-based methods in Aircraft Landing with different complexity
4. Adaptability Study
Similar Environments. In our evaluation, we intentionally introduced scenarios where two environments were very similar, with differences limited to just one grid in the grid world setting. The agent was then subjected to the unlearning process in one of these environments. After unlearning, we carefully examined the agent’s performance in the other environment.
Refer to caption (a) Steps of the four methods Refer to caption (b) Rewards of the four methods
Figure 28. The four methods in two similar environments, denoted as $A$ and $B$, in the grid world setting
The results of this experiment, shown in Figure 28, provide valuable insights. They demonstrate that when an agent undergoes unlearning in one of two similar environments, it does not exhibit deteriorated performance in the other environment. This finding supports the idea that the agent’s unlearning is environment-specific and does not extend to other environments, even if they are almost identical. This can be attributed to the characteristics of the two proposed unlearning methods. In the decremental reinforcement learning-based method, the agent is fine-tuned in the unlearning environment using a new loss function. In contrast, the poisoning-based method involves the agent’s retraining in a modified version of the unlearning environment. Consequently, even though the two environments are almost identical, the agent undergoes different training experiences, leading to distinct performance outcomes in the two environments. This outcome underscores the effectiveness of our reinforcement unlearning methods in achieving environment-specific knowledge removal while maintaining performance in other environments.
We have also undertaken the challenge of unlearning in both of the two similar environments. The outcomes, illustrated in Figure 29, reveal a deterioration in the agent’s performance across both environments. This observation underscores the profound impact of the unlearning process on the agent’s adaptability and proficiency in similar yet distinct environments. The discernible decline in performance serves as a compelling testament to the intricacies involved in unlearning an agent, shedding light on the nuanced dynamics that govern its responses in comparable scenarios.
Refer to caption (a) Steps of the four methods Refer to caption (b) Rewards of the four methods
Figure 29. The four methods in two similar environments, denoted as $A$ and $B$, in the grid world setting, when both environments are unlearned
We further investigated the scenario in which two environments are completely identical. However, the results revealed an intriguing outcome: after the agent undergoes the unlearning process in one environment, its performance in the identical counterpart deteriorates. Our hypothesis is that the agent’s updated knowledge, acquired during unlearning in one environment, is seamlessly applied to the identical environment, leading to a decline in performance. We leave the differentiation between two identical environments during reinforcement unlearning as our future research.
Robustness. The results observed in the aircraft landing (Figure 30), virtual home (Figure 31) and maze explorer (Figure 32) scenarios closely align with those in the grid world setting, showcasing the robustness of both the decremental reinforcement learning-based and poisoning-based methods against external noise. In the remaining three scenarios, the agent’s behavior continues to exhibit consistent patterns even when external noise is introduced during the unlearning process. This consistency is crucial for practical real-world applications, where agents must maintain their adaptability and performance despite uncertainties and disturbances. For example, in autonomous driving, noise from sensor readings or unexpected environmental conditions may often be encountered. The ability of the unlearning methods to maintain their efficacy despite such disturbances reinforces their practicality and reliability in dynamic environments.
Refer to caption (a) Average steps of the decremental RL-based method Refer to caption (b) Rewards of the decremental RL-based method Refer to caption (c) Average steps of the poisoning-based method Refer to caption (d) Rewards of the poisoning-based method
Figure 30. The decremental RL-based and poisoning-based methods in Aircraft Landing with and without noise
Refer to caption (a) Average steps of the decremental RL-based method Refer to caption (b) Rewards of the decremental RL-based method Refer to caption (c) Average steps of the poisoning-based method Refer to caption (d) Rewards of the poisoning-based method
Figure 31. The decremental RL-based and poisoning-based methods in Virtual Home with and without noise
Refer to caption (a) Average steps of the decremental RL-based method Refer to caption (b) Rewards of the decremental RL-based method Refer to caption (c) Average steps of the poisoning-based method Refer to caption (d) Rewards of the poisoning-based method
Figure 32. The decremental RL-based and poisoning-based methods in Maze Explorer with and without noise
5. Numerical Results
In these numerical results, a new evaluation metric is adopted: trajectory similarity, which is defined as a sequence of state-action pairs. We evaluate the effectiveness of unlearning by comparing the similarity between two trajectories: one captured before unlearning and the other obtained after unlearning. Consider two trajectories, $\tau=((s_{1},a_{1}),\ldots,(s_{m},a_{m}))$ and $\tau^{\prime}=((s^{\prime}_{1},a^{\prime}_{1}),\ldots,(s^{\prime}_{m^{\prime}},a^{\prime}_{m^{\prime}}))$, where $\tau$ represents the trajectory recorded before unlearning, and $\tau^{\prime}$ corresponds to the trajectory collected after unlearning. The similarity between $\tau$ and $\tau^{\prime}$ is computed as:
(12) $sim(\tau,\tau^{\prime})=\frac{\sum^{m}_{i=1}\mathbbm{1}_{s_{i}=s^{\prime}_{j}\wedge a_{i}=a^{\prime}_{j}}}{m},$
where $\mathbbm{1}$ is the indicator function and $j$ is an index in $[1,m^{\prime}]$.
Table 1 numerically showcases the performance of the decremental RL-based method in the grid world setting with the size of $10\times 10$. An intriguing observation is the notable difference in trajectory similarity between the unlearning environment and the retained environments. In the unlearning environment, the trajectory similarity is as low as $35\%$, indicating that the unlearned policy operates substantially differently from the original policy in this specific environment. However, in the retained environments, the trajectory similarity remains around $85\%$ in average, implying that the unlearned policy only exhibits slight variations compared to the original policy. These results support the notion of successful unlearning, as the retained environments maintain a high level of performance similarity.
Table 1. Grid world $10\times 10$ under the decremental RL-based method, where Environment 1 is the unlearning environment
Steps pre-unlearning Steps post-unlearning Reward pre-unlearning Reward post-unlearning Trajectory similarity
Environment 1 19.34 55.8 -44 -273.5 35%
Environment 2 $20.08$ $22.38$ $-47.8$ $-72.5$ $89\%$
Environment 3 $19.36$ $23.7$ $-44.6$ $-78.7$ $79\%$
Environment 4 $19.4$ $20.8$ $-45.1$ $-51.9$ $91\%$
Environment 5 $21.08$ $20.44$ $-62.2$ $-50.8$ $88\%$
Environment 6 $20.44$ $21.22$ $-50.8$ $-55.5$ $93\%$
Environment 7 $19.54$ $20.58$ $-45.9$ $-51.1$ $94\%$
Environment 8 $19.48$ $20.52$ $-44.6$ $-50.8$ $94\%$
Environment 9 $22.36$ $20.36$ $-66.8$ $-48.8$ $75\%$
Environment 10 $19.38$ $20.94$ $-46$ $-53$ $92\%$
Environment 11 $20.08$ $20.82$ $-48.5$ $-52.6$ $94\%$
Environment 12 $19.42$ $21.12$ $-45$ $-54.4$ $89\%$
Environment 13 $19.86$ $21.02$ $-47$ $-52.1$ $94\%$
Environment 14 $20.04$ $21.98$ $-48.9$ $-65.7$ $89\%$
Environment 15 $19.96$ $20.78$ $-48$ $-52.3$ $93\%$
Environment 16 $19.98$ $21.46$ $-47.3$ $-57.3$ $92\%$
Environment 17 $25.78$ $22.34$ $-110.5$ $-65.6$ $70\%$
Environment 18 $19.36$ $26.7$ $-44.7$ $-110.9$ $72\%$
Environment 19 $19.6$ $26.7$ $-45.9$ $-110.9$ $72\%$
Environment 20 $19.34$ $20.1$ $-44.1$ $-48.5$ $95\%$
The observation in the aircraft landing setting in Table 2 is similar to the grid world setting. The unlearning process successfully reduces the trajectory similarity in the unlearning environments, with values as low as $28\%$, while maintaining relatively high trajectory similarity around $90\%$ in average in the retained environments. These consistent results provide strong evidence of the effectiveness of our unlearning method in selectively modifying the agent’s behavior within the targeted unlearning environments while preserving its performance in other retained environments.
The results of the poisoning-based method align closely with those of the decremental RL-based method in the virtual home and maze explorer settings shown in Tables 3 and 4, respectively. The trajectory similarity is as low as $46\%$ and $35\%$ in the unlearning environment while maintaining high around $87\%$ and $84\%$ in average in the remaining environments in virtual home and maze explorer, respectively. This similarity can be attributed to the effectiveness of both methods in modifying the agent’s behavior within the targeted unlearning environments while preserving its performance in other retained environments. The poisoning-based method introduces deliberate changes to the state transition function of the unlearning environment, leading the agent to learn new, albeit incorrect, knowledge specific to the modified environment. Similarly, the decremental RL-based method selectively forgets learned knowledge related to the unlearning environment through iterative adjustments to the agent’s policy. As a result, both methods successfully enable the agent to adapt to the unlearning environment while maintaining its performance in other retained environments. The consistent performance across different settings showcases the robustness and efficacy of our reinforcement unlearning methods.
Table 2. Aircraft landing $10\times 10$ under the decremental RL-based method, where Environment 1 is the unlearning environment
Steps pre-unlearning Steps post-unlearning Reward pre-unlearning Reward post-unlearning Trajectory similarity
Environment 1 10.72 36.9 50 -115 28%
Environment 2 $11.4$ $12.14$ $46$ $42.3$ $91\%$
Environment 3 $11.88$ $13.38$ $43.7$ $35.6$ $89\%$
Environment 4 $11.52$ $12.42$ $45.4$ $40.7$ $90\%$
Environment 5 $10.86$ $12.36$ $49.2$ $39.6$ $84\%$
Environment 6 $10.1$ $11.88$ $53.2$ $42.7$ $79\%$
Environment 7 $10$ $12.08$ $53$ $42.8$ $81\%$
Environment 8 $10.02$ $11.6$ $53.1$ $44.8$ $82\%$
Environment 9 $11.62$ $13.2$ $44.5$ $36.6$ $88\%$
Environment 10 $10.02$ $11.84$ $53.9$ $43$ $83\%$
Environment 11 $10.12$ $12.12$ $53.2$ $42.6$ $82\%$
Environment 12 $11$ $12.48$ $48.3$ $40.4$ $82\%$
Environment 13 $12.06$ $13.34$ $42.6$ $35.5$ $89\%$
Environment 14 $10.44$ $11.4$ $50.6$ $46.2$ $89\%$
Environment 15 $11.1$ $11.9$ $47.7$ $43.6$ $91\%$
Environment 16 $10.28$ $12.18$ $52.3$ $41.8$ $82\%$
Environment 17 $9.94$ $11.72$ $53.8$ $43$ $85\%$
Environment 18 $11.1$ $12.72$ $48$ $39.3$ $83\%$
Environment 19 $10.08$ $12.12$ $53.2$ $42.3$ $78\%$
Environment 20 $11.4$ $11.74$ $46.2$ $44.8$ $95\%$
Table 3. Virtual home under the poisoning-based method, where Environment 1 is the unlearning environment
Steps pre-unlearning Steps post-unlearning Reward pre-unlearning Reward post-unlearning Trajectory similarity
Environment 1 26.34 55.6 92.88 6.48 46%
Environment 2 $23.42$ $25.02$ $102$ $94.3$ $86\%$
Environment 3 $23.76$ $24.02$ $90.56$ $87.36$ $94\%$
Environment 4 $28.28$ $30.48$ $80.44$ $82.48$ $89\%$
Environment 5 $31.5$ $32.3$ $76.28$ $72.6$ $90\%$
Environment 6 $33.38$ $35.04$ $74.2$ $70.6$ $93\%$
Environment 7 $27.06$ $26.96$ $89.56$ $92.16$ $99\%$
Environment 8 $27.84$ $31.64$ $91.28$ $80.72$ $85\%$
Environment 9 $21.28$ $22.8$ $100.16$ $94.16$ $87\%$
Environment 10 $32.28$ $32.2$ $80.88$ $81.8$ $96\%$
Table 4. Maze explorer under the poisoning-based method, where Environment 1 is the unlearning environment
Steps pre-unlearning Steps post-unlearning Reward pre-unlearning Reward post-unlearning Trajectory similarity
Environment 1 35.2 77.3 88 21.2 35%
Environment 2 $37.2$ $37.9$ $83.6$ $83.4$ $86\%$
Environment 3 $34.9$ $36.1$ $87.51$ $85.47$ $80\%$
Environment 4 $33.1$ $33.9$ $91.6$ $90.1$ $84\%$
Environment 5 $35.9$ $37.7$ $86.81$ $81.75$ $80\%$
Environment 6 $36.4$ $36.3$ $83.9$ $85.1$ $90\%$
Environment 7 $38.9$ $40.1$ $81.7$ $79.67$ $87\%$
Environment 8 $35.7$ $38.1$ $87.15$ $81.1$ $79\%$
Environment 9 $34.5$ $35.5$ $89.2$ $86.5$ $82\%$
Environment 10 $36.1$ $37.2$ $85.47$ $82.6$ $85\%$

### v5

@@ VII Conclusion
This paper presents a pioneering research area, termed reinforcement unlearning, which addresses the crucial need to protect the privacy of environment owners by enabling an agent to unlearn entire environments. We propose two distinct reinforcement unlearning methods: decremental RL-based and environment poisoning-based approaches. These methods are designed to be adaptable to different situations and provide effective mechanisms for unlearning. Also, we introduce a novel concept termed “environment inference” to evaluate the outcomes of the unlearning process.
Our future work focuses on developing a unified framework for reinforcement unlearning that integrates various unlearning methods and supports easy application to different RL algorithms and environments. This includes designing a modular architecture that allows for interchangeable unlearning components tailored to specific needs of different RL scenarios.
Acknowledgments
Tianqing Zhu is the corresponding author. This research is supported by the Australian Research Council (ARC) Discovery Grant DP230100246. Moreover, Minhui Xue’s work is partially supported by ARC Grant DP240103068 and by the CSIRO–National Science Foundation (US) AI Research Collaboration Program.
@@ Acknowledgments
I Unlearning in a Single Environment
Problem Definition. Unlearning within a single environment primarily focuses on selectively forgetting specific knowledge from that environment. We formally define the problem as follows. Given an environment =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{T},r\}$, a set of $n$ trajectories $(\tau_{1},...,\tau_{n})$ occur within it. Each trajectory $\tau_{i}=((s^{i}_{1},a^{i}_{1}),\ldots,(s^{i}_{k},a^{i}_{k}))$ is a sequence of $k$ states and actions experienced by the agent as it interacts with the environment according to a policy. Assume the agent aims to unlearn a trajectory $\tau_{u}$, the goal is to update the policy $\pi$ to $\pi^{\prime}$ such that the reward obtained along trajectory $\tau_{u}$ is minimized:
$\min_{\pi^{\prime}}\mathbb{E}_{(s,a)\sim\tau_{u}}[Q_{\pi^{\prime}}(s,a)],$ (12)
while unaffecting the reward in other trajectories:
$\min_{\pi^{\prime}}\mathbb{E}_{(s,a)\not\sim\tau_{u}}|Q_{\pi^{\prime}}(s,a)-Q_{\pi}(s,a)|.$ (13)
In Eq. 12, minimizing the reward for specific state-action pairs teaches the agent these pairs are less valuable. Consequently, the agent may select alternative actions in those states to avoid the reduced rewards. This change in action selection leads to different subsequent states, effectively altering the unlearned trajectory. Essentially, this makes the trajectory unlikely to occur, achieving the effect of forgetting it. In comparison, Eq. 13 focuses on preserving the integrity of the remaining trajectories, effectively ensuring they remain unaffected. This goal is to maintain stability and reliability in the agent’s performance across these trajectories, emphasizing remembering them, while selectively forgetting the specified trajectory.
Methods. To adapt our methods to unlearning in a single-environment scenario, we have modified the loss and reward functions accordingly. For the decremental RL-based method, the revised loss function is:
$\mathcal{L}_{u}=\mathbb{E}_{(s,a)\sim\tau_{u}}[Q_{\pi^{\prime}}(s,a)]+\mathbb{E}_{(s,a)\not\sim\tau_{u}}|Q_{\pi^{\prime}}(s,a)-Q_{\pi}(s,a)|,$
where the first term reduces rewards along the trajectory $\tau_{u}$, and the second term maintains policy performance elsewhere. Similarly, the reward function for the poisoning-based method is reformulated as:
$\mathcal{R}_{i}:=\lambda_{1}\Delta(\pi_{i}(s_{i})||\pi_{i+1}(s_{i}))+\lambda_{2}\sum_{s\not\sim\tau_{u}}\sum_{a}\pi_{i}(s,a)r(s,a),$
targeting a higher divergence between successive policies $\pi_{i}$ and $\pi_{i+1}$ and ensuring good performance on remaining trajectories.
Experimental Results. The results of unlearning within a single environment in the grid world scenario, depicted in Figure 25, align with those observed in multi-environment settings: the agent performs deterioratively in the unlearning trajectory, receiving lower rewards, while continuing to perform well in other trajectories. Note that counting steps is less relevant as an evaluation metric in the single-environment setting compared to multi-environment settings, because the focus is now on ensuring the agent effectively avoids specific states or actions, rather than on the speed or distance covered.
(a) Rewards of decremental RL-based
(b) Rewards of poisoning-based
Figure 25: The two methods in unlearning Trajectory 1 within a single environment using the grid world setting.
II Model Architecture
In the grid world and aircraft landing settings, we employ a fully-connected neural network as the model for our reinforcement learning agent. The neural network architecture consists of an input layer, two hidden layers, and an output layer. The input layer takes a 10-dimensional vector as input, representing the relevant features of the environment. The output layer generates a 4-dimensional vector, representing the probability distribution over the four possible actions: up, down, left, and right. The first hidden layer comprises 64 neurons, while the second one consists of 32 neurons. Similarly, for recommendation systems, the architecture is slightly adjusted with an input layer, two hidden layers, and an output layer with neuron counts of 4, 128, 128, and 2 respectively. The input layer manages four movie attributes: ID, release year, genre, and rating, while the output layer determines whether to recommend the movie.
@@ II Model Architecture
III Comparison of Various RL Algorithms
(a) The average number of steps with different RL algorithms
(b) The average rewards received by the agent with different RL algorithms
Figure 26: Performance of the three RL algorithms in grid world.
To assess the effectiveness of different RL algorithms in our setting, we evaluate the performance of PPO and DDPG and compare them with the adopted DQN algorithm. As illustrated in Figure 26, PPO and DDPG closely mirrors that of the DQN algorithm during the unlearning process, albeit with inferior results. Specifically, the agent achieves lower rewards and requires more steps to complete tasks. Several factors may contribute to this disparity. Firstly, DQN’s exploration-exploitation trade-off may be better suited for efficiently navigating complex environments. Additionally, its experience replay mechanism enhances sample efficiency and learning convergence, further bolstering its performance. Thus, PPO and DDPG are not extensively utilized in our experiments.